OpenStudy (anonymous):

The function H described by H(x)= 2.75+71.48 can be used to predict the height, in centimeters, of a woman whose humurus (the bone from the elbow to the shoulder) is x cm long. Predict the height of the woman whose humerus is 34 cm long. HEY SAIFOO...Yes, it is me!

OpenStudy (anonymous):

you are still missing an x!

OpenStudy (saifoo.khan):

I knw the answer! :O

OpenStudy (anonymous):

\[H(x)= 2.75+71.48\]

OpenStudy (saifoo.khan):

\[H(x)= 2.75x+71.48 \]

OpenStudy (anonymous):

is it \[H(x)= 2.75+71.48x\] or \[H(x)= 2.75x+71.48\]

OpenStudy (saifoo.khan):

i did this problem before!

OpenStudy (anonymous):

first karnak award of the day

OpenStudy (saifoo.khan):

\[H(x)= 2.75(34)+71.48\]

OpenStudy (saifoo.khan):

whts this? @satallite.

OpenStudy (saifoo.khan):

164.98

OpenStudy (anonymous):

something old http://www.youtube.com/watch?v=XnwyQFe3wRA

OpenStudy (anonymous):

That's satellite

OpenStudy (anonymous):

he was given answers, found the questions

OpenStudy (saifoo.khan):

Haha.

OpenStudy (saifoo.khan):

everytime he posts the same thing!

OpenStudy (anonymous):

that is the karnak award given for questions that say things like "pleez help f(x)=2x+5"

OpenStudy (saifoo.khan):

Haha.

OpenStudy (saifoo.khan):

whoz Karnak?

OpenStudy (anonymous):

or "here is my problem 2x+3"

OpenStudy (anonymous):

old johnny carson sketch character

OpenStudy (saifoo.khan):

O k.

OpenStudy (saifoo.khan):

So dumb? u got the answer?

OpenStudy (anonymous):

Thank you Saifoo!!!

OpenStudy (anonymous):

I am clicking on good answer but it is not doing anything!

OpenStudy (saifoo.khan):

refresh the page! ^^